what is all silver tea cups in trignonmetery i cant understand its application and                                                                                                what is value of alpha in tan alpha=cot a/2

Grade:10

1 Answers

maurya ksk
11 years ago

here ispecify alpla as a

  Given,

    tan a=cot a/2

cot(90-a)=cot a/2

so,

     90-a=a/2

     90=a+a/2

     90=3a/2

     a=90*2/3

     a=60